OpenStudy (anonymous):

The standard golf ball has a mass of 45g. How many golf balls are there in a bag weighing 1.8kg?

OpenStudy (anonymous):

Is it 400?

OpenStudy (paxpolaris):

\[1.8kg \times{1000 g \over 1kg} \times {1 ball \over 45g}\]

OpenStudy (anonymous):

wait no is it 40?

OpenStudy (paxpolaris):

yes 40 balls

OpenStudy (mayankdevnani):

\[\large \bf 1~ball~weigh~45g=0.045kg\] so, \[\large \bf 1kg~weigh=\frac{1}{0.045}~balls\] but we have to find out,1.8 kg,so \[\large \bf 1.8kg~weigh=\frac{1}{0.045} \times 1.8=40~balls=\color{red}{ANSWER}\]

OpenStudy (mayankdevnani):

hope you understand. @Binkyboo100

OpenStudy (anonymous):

thanks I do now @mayankdevnani

OpenStudy (mayankdevnani):

good :)